OpenStudy (woznichak):

Please help Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

rishavraj (rishavraj):

so u see g(-1) = -1 + 3 g(-1) = 2 now u need to find f(g(x))... and g(-1) = 2 so f(2) = 5*2 - 10 = 0

OpenStudy (woznichak):

So basically you just fill it in and solve it then use the first equation to solve the second one?

rishavraj (rishavraj):

see it says f(g(x)) at x = -1 so first u need to calculate g(x) at x = -1